4. You have a solution with the same concentration of two proteins that bind glucose with different affinities GB1, with a Kd of 10 mM, and GB2, with a Kd of 100mM. i) Which protein has higher affinity for glucose? i) what will be the % occupancies of GB1 and GB2 at glucose concentrations of 0.05ml, 0.1 mM. 0.5mM, and 1mM?


Answer part ii.
Answer #1

Kd is also known as disassociation constant. It is defined as the concentration at which 50% of the binding sites (receptors) are occupied by the drug.

a) The disassociation constant is inversely proportional to the affinity of substrate. Therefore, GB1 which has lower Kd has the higher affinity for glucose.

b) Percent occupancies of GB1:

At 0.05 mM concentration,

= 0.05/(0.05+10)

= 0.05/10.05

= 0.49%

At 0.1mM concentration,

= 0.1/(0.1+10)

= 0.99%

At 0.5 mM concentration,

= (0.5)/(0.5+10)

= 4.76%

At 1 mM concentration,

= 1/(1+10)

= 9.1%

Percent occupancies of GB2:

At 0.05 mM concentration,

= 0.05/(0.05+100)

= 0.05%

At 0.1 mM concentration,

= 0.1/(0.1+100)

= 0.1%

At 0.5 mM concentration,

= 0.5/(0.5+100)

= 0.49%

At 1 mM concentration,

= 1/(1+100)

= 0.99%